OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -4 and a5 = -1,024?

OpenStudy (tkhunny):

a1 = -4 a2 = -4r a3 = -4r^2 a4 = -4r^3 a5 = -4r^4 = -1024 a6 = -4r^5 a7 = -4r^6 I'm seeing a pattern.